Cottonwood Company reports the following operating results for the month of April.

COTTONWOOD COMPANY
CVP Income Statement
For the Month Ended April 30, 2020
Total Per Unit
Sales (8,500 units) $382,500 $45
Variable costs 153,000 18
Contribution margin 229,500 $27
Fixed costs 178,200
Operating income $51,300


Management is considering the following course of action to increase operating income: Reduce the selling price by 20%, with no changes to unit variable costs or fixed costs. Management is confident that this change will increase unit sales by 30%.

Using the contribution margin technique, calculate the break-even point in units and dollars and margin of safety in dollars, assuming changes to sales price and volume as described above.
Answer #1

Solution:

New selling price per unit = $45 *80% = $36

New contribution margin per unit = $36 - $18 = $18 per unit

CM ratio = $18/$36 = 50%

Fixed costs = $178,200

Break-even point in units = Fixed costs / CM per unit = $178,200 / $18 =9900 units

Break-even point in dollars = Fixed costs / CM ratio = $178,200 / 50% = $356,400

New sales volume = 8500*130% = 11050 units

New sales revenue = 11050*$36 = $397,800

Margin of safety in dollars = Current sales - Breakeven sales

= $397,800 - $356,400 = $41,400